CHARACTERIZATION OF C_{3}HSi^{−}

\underline{N. INOSTROZA}, Departamento de Astrofísica Molecular e Infrarroja, Instituto de Estructura de la Materia, C.S.I.C., Serrano 121, Madrid 28006, SPAIN; M.L. SENENT, Departamento de Astrofísica Molecular e Infrarroja, Instituto de Estructura de la Materia, C.S.I.C., Serrano 121, Madrid 28006, SPAIN.

Structural and spectroscopic properties of the astrophysically relevant SiC_{3}H^{−} anion are investigated at the RCCSD(T) and CASSCF levels of theory. The molecule presents a total of 12 isomers (4 lineal and 8 cyclic), all of them showing singlet ground electronic states. The stability of the linear isomer l − SiC_{3}H^{−}(X^{1}Σ^{−}) is notorious. The B0 rotational constant of l − Si^{12}C_{3}H^{−} and l − Si^{12}C_{3}D^{−}, the dipole moment, excitation energies to the lowest electronic states and electron affinities have been estimated.